# CollabLLM: From Passive Responders to Active Collaborators

### v1

@@ 1 Introduction
Here we introduce CollabLLM, a novel and general training framework that improves the ability of LLMs to effectively collaborate with humans in multiturn scenarios (Gao et al., 2019; Balog & Zhai, 2023; Rahmani et al., 2023). The key innovation of CollabLLM is to promote LLMs’ forward-looking behavior that leads to long-term collaboration gains (Figure 1). We introduce a collaborative simulation module that samples future conversations with users to estimate the long-term impact of model responses across multiple turns, a measure we term the Multiturn-aware Reward (MR). The MR function evaluates responses by incorporating both extrinsic metrics, such as task-specific success, and intrinsic metrics, such as efficiency, to holistically assess collaboration quality (cf. Section 3). By fine-tuning with established RL algorithms (Rafailov et al., 2023; Schulman et al., 2017) on MRs, CollabLLM promotes responses that lead to better task completion and efficiency in later conversation stages. As shown in Figure 2b, the fine-tuned model goes beyond simply responding to user requests in Figure 2a—it actively collaborates by asking follow-up questions about the writing tone, generating targeted content about the role of optimism, and offering insightful suggestions such as adding anecdotes.
@@ 3.1.1 Conversation-level Reward Function
where focuses on task-specific success, and evaluates user experience including efficiency and engagement.
@@ 3.2 Optimization & Synthetic Datasets
4 Experimental Setup***Dataset and training details in Appendix B; all prompts (e.g., prompts of user simulator and LLM judges) in Appendix C.
@@
In addition to the above task-specific metrics, we incorporate two task-agnostic scores across all datasets: 1) Average Token Count, which quantifies the average number of tokens generated by the LLM per conversation, reflecting interaction efficiency. 2) Interactivity (ITR), which evaluates engagement levels using an LLM judge (Claude-3.5-Sonnet), with scores rescaled to the range [0, 1].
@@ 5 Results of Simulated Experiments
CollabLLM increases task performance, conversational efficiency, and engagement. CollabLLM achieves 18.5% superior task-specific performance, 13.3% more efficient conversations, and 46.3% enhanced interactivity compared to the best baselines. We highlight that CollabLLM engage in more meaningful collaborations, with ITR shows substantial gains. For MediumDocEdit-Chat, the Online DPO model increases ITR from 0.46 to 0.92. Moreover, our framework significantly improves conversational efficiency by minimizing the content users need to review to arrive at the final solution. For MATH-Chat, Online DPO decreases token count per conversation by 1.03k compared to the base model.
@@ 5.3 Model Generalization (Table 2)
We assess CollabLLM, trained with online DPO on BigCodeBench-Chat (the coding assistance task), on Abg-CoQA (Guo et al., 2021), a question-answering (QA) benchmark where questions are labeled as ambiguous or non-ambiguous (cf. Appendix D). We categorize the model’s responses into two actions—asking a clarifying question or providing a direct answer—and evaluate action-level accuracy within each question type. As shown in Table 2, GPT-4o and Llama-3.1-8B rarely ask clarifying questions regardless of ambiguity. In contrast, CollabLLM proactively asks questions about 50% of the time while maintaining high accuracy on unambiguous inputs. This behavior leads to the highest Macro Accuracy across both ambiguous and non-ambiguous sets and improves Macro F1 over the base model, while leaving room for further improvement against GPT-4o. These results suggest that CollabLLM effectively generalizes its learned collaborative strategies beyond its training domain.
@@ 6 Real-world User Study
Setup. We conduct a large-scale user study using Amazon Mechanical Turk with 201 participants. Each participant is assigned a document type—randomly selected to be either blog post, creative writing, or personal statement—and chooses a topic from a predefined set. To simulate real-world scenarios where users have only a rough idea of the task, they are first asked to provide brief responses to topic-related questions. Participants then engage in at least eight turns of conversation with an anonymized AI assistant, which can be Base, Proactive Base, or CollabLLM. Every three turns, they provide an interaction rating based on their experience so far. After the conversation, participants rate the final document quality and overall interaction. All ratings are in a scale from 1 to 10. We also record the total interaction duration to assess efficiency. The detailed user study setup is provided in Appendix E.
@@ 7 Related Work
Non-collaborative LLM training. Existing LLM training frameworks, including pre-training, supervised fine-tuning (SFT), and reinforcement learning (RL) (Rafailov et al., 2023; Schulman et al., 2017; Ouyang et al., 2022; Lee et al., 2024), primarily optimize for next-turn response quality. Standard RL methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) apply rewards to individual model responses without accounting for their long-term impact on conversation trajectories. While effective for single-turn objectives, these approaches fail to capture how responses influence user intent discovery and long-term task success (Amershi et al., 2019; Zamfirescu-Pereira et al., 2023; Wang et al., 2024; Kim et al., 2024; Lu et al., 2024).
@@
LLMs for generating clarification questions: Beyond prompting, prior studies have explored supervised fine-tuning (Andukuri et al., 2024), RL fine-tuning (Chen et al., 2024; Zamani et al., 2020; Erbacher & Soulier, 2023), and active learning (Pang et al., 2024) to train models to ask clarification questions. For example, Chen et al. (2024) use Direct Preference Optimization (DPO) to encourage models to identify ambiguous user queries and request clarifications. However, like prompting approaches, these methods primarily focus on clarification questions and do not generalize to broader multiturn collaboration strategies.
Multiturn training for LLMs: Several studies extend RLHF to multiturn settings by optimizing trajectory-level rewards (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024; Shi et al., 2024b). Other works (Xu et al., 2023; Deng et al., 2024) leverage self-chat or self-play to enhance model adaptation to dialogue. See Zhang et al. (2025) for a comprehensive survey. However, these methods primarily rely on post-hoc trajectory-level data, learning from observed conversations rather than explicitly modeling the causal effect of individual responses on long-term task success (see Appendix A for further details). Additionally, they often overlook open-ended tasks such as document generation (Faltings et al., 2023; Jiang et al., 2024), where real-world user responses can be highly variable, and users may have limited capacity to read and refine lengthy model outputs.
@@ 8 Conclusion
Multiturn human-LLM collaborations are increasingly prevalent in real-world applications. We believe future foundation models should act as collaborators rather than passive responders, actively uncovering user intents in open-ended and complex tasks—an area where current LLMs fall short. The key insight of CollabLLM is making LLMs more multiturn-aware by using forward sampling to estimate the long-term impact of responses. Through extensive simulated and real-world evaluations, we demonstrate that CollabLLM is highly effective, efficient, and engaging, while also generalizing well to new tasks and interactions, advancing the frontiers of human-centered LLMs.
@@
This work presents one of the first attempts to train LLMs in such human-centric environments. To promote future research in this societally beneficial direction, we plan to release all the code, models, data, benchmarks, and user simulators described in this work.
@@
Acknowledgments
We thank Vishal Chowdhary, Hoifung Poon, Swadheen Shukla, Chandan Singh and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions. We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback.
@@ A.2 Distinction from Other Multiturn Training Frameworks
Existing multiturn trajectory-based training frameworks (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024) primarily rely on learning from observed trajectory-level rewards. These methods estimate the utility of responses by assigning rewards post hoc to completed conversations, training models to imitate historically successful responses. However, this approach is fundamentally correlational—it captures statistical associations between responses and final outcomes but does not explicitly model how different responses causally influence future turns.
In contrast, our Multiturn-aware Reward (MR) framework intervenes on model responses and uses forward sampling to explicitly simulate alternative future trajectories. This enables the model to estimate the counterfactual impact of different responses and adjust its behavior accordingly. By leveraging causal effect estimation, MR training moves beyond passive imitation of high-reward conversations and instead actively optimizes responses to maximize long-term goal attainment. This distinction is crucial in dynamic human-LLM interactions where user needs evolve throughout a conversation.
Appendix B Dataset and Training Details
@@ B.2 Training Details
We provide the hyperparameters for CollabLLM fine-tuning in Table 6.
@@
Appendix C Prompts
C.1 User Simulator
C.2 Prompt for Proactive Base
C.3 System Prompt
C.4 Interactivity Metric by LLM Judge
C.5 Helpfulness Reward by LLM Judge
Appendix D Question Template and Example on Abg-CoQA
@@
Appendix E User Study

### v2

@@ 1 Introduction
Here we introduce CollabLLM, a novel and general training framework that improves the ability of LLMs to effectively collaborate with humans in multiturn scenarios (Gao et al., 2019; Balog & Zhai, 2023; Rahmani et al., 2023). The key innovation of CollabLLM is to promote LLMs’ forward-looking behavior that leads to long-term collaboration gains (Figure 1). We introduce a collaborative simulation module that samples future conversations with users to estimate the long-term impact of model responses across multiple turns, a measure we term the Multiturn-aware Reward (MR). The MR function evaluates responses by incorporating both extrinsic metrics, such as task-specific success, and intrinsic metrics, such as efficiency, to holistically assess collaboration quality (cf. Section 3). By fine-tuning with RL algorithms (Rafailov et al., 2023; Schulman et al., 2017) on MRs, CollabLLM promotes responses that lead to better task completion and efficiency in later conversation stages. As shown in Figure 2b, the fine-tuned model goes beyond simply responding to user requests in Figure 2a—it actively collaborates by asking follow-up questions about the writing tone, generating targeted content about the role of optimism, and offering insightful suggestions such as adding anecdotes.
@@ 3.1.1 Conversation-level Reward Function
where focuses on task success, and evaluates user experience including efficiency and engagement.
@@ 3.2 Optimization & Synthetic Datasets
4 Experimental Setup***Dataset and training details in Appendix B; all prompts (e.g., prompts of user simulator and LLM judges) in Appendix D.
@@
In addition to the above task-specific metrics, we incorporate two task-agnostic scores across all datasets: 1) Average Token Count, which quantifies the average number of tokens generated by the LLM per conversation, reflecting interaction efficiency. 2) Interactivity (ITR), which evaluates engagement levels using an LLM judge (Claude-3.5-Sonnet), with scores rescaled to an upper bound of 1.
@@ 5 Results of Simulated Experiments
CollabLLM improves task performance, efficiency, and engagement. CollabLLM achieves 18.5% superior task-specific performance, 13.3% more efficient conversations, and 46.3% enhanced interactivity compared to the best baselines. We highlight that CollabLLM engage in more meaningful collaborations, with ITR shows substantial gains. For MediumDocEdit-Chat, the Online DPO model increases ITR from 0.46 to 0.92. Moreover, our framework significantly improves conversational efficiency by minimizing the content users need to review to arrive at the final solution. For MATH-Chat, Online DPO decreases token count per conversation by 1.03k compared to the base model.
@@ 5.3 Model Generalization (Table 2)
We assess CollabLLM, trained with online DPO on BigCodeBench-Chat (the coding assistance task), on Abg-CoQA (Guo et al., 2021), a question-answering (QA) benchmark where questions are labeled as ambiguous or non-ambiguous (cf. Appendix E). We categorize the model’s responses into two actions—asking a clarifying question or providing a direct answer—and evaluate action-level accuracy within each question type. As shown in Table 2, GPT-4o and Llama-3.1-8B rarely ask clarifying questions regardless of ambiguity. In contrast, CollabLLM proactively asks questions about 50% of the time while maintaining high accuracy on unambiguous inputs. This behavior leads to the highest Macro Accuracy across both ambiguous and non-ambiguous sets and improves Macro F1 over the base model, while leaving room for further improvement against GPT-4o. These results suggest that CollabLLM effectively generalizes its learned collaborative strategies beyond its training domain.
@@ 6 Real-world User Study
Setup. We conduct a large-scale user study using Amazon Mechanical Turk with 201 participants. Each participant is assigned a document type—randomly selected to be either blog post, creative writing, or personal statement—and chooses a topic from a predefined set. To simulate real-world scenarios where users have only a rough idea of the task, they are first asked to provide brief responses to topic-related questions. Participants then engage in at least eight turns of conversation with an anonymized AI assistant, which can be Base, Proactive Base, or CollabLLM. Every three turns, they provide an interaction rating based on their experience so far. After the conversation, participants rate the final document quality and overall interaction. All ratings are in a scale from 1 to 10. We also record the total interaction duration to assess efficiency. The detailed user study setup is provided in Appendix F.
@@ 7 Related Work
Non-collaborative LLM training. Existing LLM training frameworks, including pre-training, supervised fine-tuning (SFT), and reinforcement learning (RL) (Rafailov et al., 2023; Schulman et al., 2017; Ouyang et al., 2022; Lee et al., 2024), primarily optimize for next-turn response quality. Standard RL methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) apply rewards to individual model responses without accounting for their long-term impact on conversation trajectories. While effective for single-turn objectives, these approaches fail to capture how responses influence user intent discovery and long-term task success (Amershi et al., 2019; Zamfirescu-Pereira et al., 2023; Wang et al., 2024; Kim et al., 2024).
@@
LLMs for generating clarification questions: Beyond prompting, prior studies have explored supervised fine-tuning (Andukuri et al., 2024), RL fine-tuning (Chen et al., 2024; Zamani et al., 2020; Erbacher & Soulier, 2023), and active learning (Pang et al., 2024) to train models to ask clarification questions. For example, Chen et al. (2024) use Direct Preference Optimization (DPO) to encourage models to request clarifications. However, like prompting approaches, these methods primarily focus on clarification questions and do not generalize to broader multiturn collaboration strategies.
Multiturn training for LLMs: Recent benchmarks (Abdulhai et al., 2023; Kwan et al., 2024) evaluate LLMs’ performance in multiturn settings, measuring the goal orientation and planning capabilities of models across interactions. Several studies extend RLHF to multiturn settings by optimizing trajectory-level rewards (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024; Shi et al., 2024b; Zhang et al., 2025). Other works (Xu et al., 2023; Deng et al., 2024) leverage self-chat or self-play to enhance model adaptation. However, these methods primarily rely on post-hoc trajectory-level data, learning from observed conversations rather than explicitly modeling the causal effect of individual responses on task success (see Appendix A for further explanations). Additionally, they often overlook open-ended tasks such as document generation (Faltings et al., 2023; Jiang et al., 2024), where user responses can be highly diverse, and users may have limited capacity to read and refine lengthy model outputs.
User simulators for enhancing AI systems. Recent works employ user simulators to enhance dialogue systems (Shi et al., 2019; Tseng et al., 2021) and LLMs (Hong et al., 2023; Hu et al., 2023; Faltings et al., 2023). Recently, Hong et al. (2023) leverage LLMs to create diverse synthetic dialogues with varying user personas to train smaller dialogue models. CollabLLM differs in leveraging user simulators in forward sampling to account for long-term effect in both offline and online training.
@@ 8 Conclusion
Multiturn human-LLM collaborations are increasingly prevalent in real-world applications. Foundation models should act as collaborators rather than passive responders, actively uncovering user intents in open-ended and complex tasks—an area where current LLMs fall short. The key insight of CollabLLM is making LLMs more multiturn-aware by using forward sampling to estimate the long-term impact of responses. Through extensive simulated and real-world evaluations, we demonstrate that CollabLLM is highly effective, efficient, and engaging, while also generalizing well to new tasks and interactions, advancing the frontiers of human-centered LLMs.
Acknowledgments
We thank Doug Burger, Vishal Chowder, Jeevana Priya Inala, Hoifung Poon, Swadheen Shukla, Chandan Singh, Desney Tan and Chenglong Wang, as well as members of the Deep Learning and Health Futures groups at Microsoft Research for helpful discussions. We thank lab members in Leskovec and Zou’s labs for discussions and for providing feedback. We also gratefully acknowledge the support of NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM), IIS-2403318 (III); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
@@ Acknowledgments
Since the models in this work are trained collaboratively and aim to better align with user intent, concerns may arise regarding users with malevolent goals. However, we argue that CollabLLM can help mitigate safety risks in such cases—at least when used with LLMs that have been aligned for safety (as is the case for all models used in this work). Safety-aligned LLMs generally refuse to respond to unsafe queries, which often leads malicious users to obscure their true intentions in order to bypass safeguards. This is where our approach offers an advantage: CollabLLM often seeks to clarify user intent, creating additional opportunities to detect misuse. For example, malicious users might unintentionally reveal their actual goals, or their vagueness and refusal to disclose motivations could raise red flags—potentially providing the LLM with further cues for identifying unsafe behavior. As presented in Appendix C, we conducted various safety experiments and show that CollabLLM performs no worse than an equivalent non-collaboratively trained model in terms of safety.
@@
This work presents one of the first attempts to train LLMs in such human-centric environments. To promote future research in this societally beneficial direction, we release all the code, models, data, benchmarks, and user simulators described in this work.
@@ A.2 Distinction from Other Multiturn Training Frameworks
Existing multiturn trajectory-based training frameworks (Shani et al., 2024; Zhou et al., 2024; Gao et al., 2024) primarily rely on learning from observed trajectory-level rewards. These methods estimate the utility of responses by assigning rewards post hoc to completed conversations, typically training models to prefer higher-rated conversations over lower-rated ones. However, this approach is fundamentally observational—it captures statistical associations between responses and final outcomes, without disentangling how individual responses causally influence future turns. For example, in MTPO (Shani et al., 2024), the learning signal remains coarse-grained: rewards are assigned at the trajectory level, and the influence of specific turns within a conversation remains confounded and indirect.
In contrast, our Multiturn-aware Reward (MR) framework intervenes on individual model responses and uses forward simulation to generate alternative future trajectories. This allows the model to estimate the counterfactual impact of different responses at each turn, thereby enabling fine-grained optimization. By leveraging causal effect estimation, MR training moves beyond passive imitation of high-reward conversations and instead actively selects responses to maximize long-term task success. This interventional approach provides turn-level credit assignment that is critical in dynamic human-LLM interactions, where user needs evolve and the consequences of early decisions compound over time.
Appendix B Experimental Details
@@ B.2 Training Details
Hyperparameters (Table 6). We provide the hyperparameters for CollabLLM fine-tuning.
@@
Training Cost (Table 7). We compute average statistics over 100 future conversations on MediumDocEdit-Chat, the document editing task, which incurs the highest computational overhead among the three tasks. The table shows that even at the largest window size (), the total per-sample cost remains low, suggesting that our multi-turn training setup is financially practical. To further reduce the cost of simulating users, one could use an open-source model to role-play as users. Unfortunately, at the current stage, we find that open-source models generally perform poorly, often getting “confused” and starting to solve problems as an assistant rather than acting as a user. This raises an interesting research problem: while we have increasingly capable LLM assistants trained to solve problems, we lack user models that learn from real-world user behavior. Building better user models could be valuable for running simulations in real-world applications.
| 0.89 | 0.42 | 7.41 | 1.85 | 0.26 | 0.00174 | |||||||||||||
| 2.55 | 0.91 | 15.84 | 4.55 | 0.69 | 0.00439 | |||||||||||||
| 4.13 | 1.22 | 21.72 | 7.18 | 1.06 | 0.00685 |
B.3 Full Ablation Results
To further understand the source of performance improvements, we conduct a full ablation by training models with isolated reward signals—Helpfulness, Extrinsic Only, and Intrinsic Only—across window sizes . The resulting BLEU, token usage, and ITR scores are reported in Figure 9.
We make three key observations:
Helpfulness alone leads to marginal improvements in BLEU and ITR, but significantly increases token usage, especially at larger window sizes, suggesting verbosity rather than improved efficiency or interactivity.
Extrinsic-only reward achieves strong BLEU scores (e.g., 0.377 at ), indicating good task alignment. However, it underperforms in ITR and often generates longer responses.
Intrinsic-only reward improves ITR at (e.g., 0.74), but offers lower BLEU and comparable or slightly lower token efficiency, indicating better interactivity at the expense of task success.
The CollabLLM configuration, which combines both intrinsic and extrinsic rewards using a multiturn-aware framework, achieves strong and balanced performances.
Note that the choice of reward type (intrinsic or extrinsic) is independent of the multiturn-aware reward design. In practice, one can flexibly plug in different reward signals, which are then used to evaluate the responses’ long-term impact through forward sampling.
Appendix C Safety Evaluation
As the models in this work are collaboratively trained and designed to be more aligned with the user’s intent, concerns may arise if a user happens to have malevolent intentions. However, we note that CollabLLM models were finetuned from Llama-3.1-8B, which has been aligned for safety—so jailbreaking CollabLLM still poses a significant challenge. To determine whether collaborative training weakens the safety features inherent to a model (Llama-3.1-8B) that has undergone significant alignment steps for safety, we performed an adversarial evaluation using the Azure AI Evaluation SDK‡‡‡https://learn.microsoft.com/en-us/python/api/overview/azure/ai-evaluation-readme and prompted both the baseline and CollabLLM with various offensive queries intended to elicit unsafe responses.
Specifically, we performed the following steps:
Adversarial query selection: We used the SDK’s AdversarialSimulator to generate adversarial queries (e.g., queries encouraging the LLM to produce hateful comments). We then used the SDK’s harm evaluators (ViolenceEvaluator, SexualEvaluator, SelfHarmEvaluator, HateUnfairnessEvaluator) to categorize each query into one of four harm types: violence, sexual, self-harm, and hate. For each query, we used the highest score among the four evaluators to determine its harm category. We randomly selected 20 adversarial queries per harm category, resulting in a total of 80 queries.
Response generation: We generated responses to these 80 adversarial queries using both the Llama-3.1-8B baseline model and CollabLLM.
Harm scoring: We evaluated each model-generated response using all four harm evaluators to ensure comprehensive assessment.
| Model | Harm score (0–7 range, ) | |||
| Violence | Sexual | Self-harm | Hate | |
| Llama-3.1-8B | 0.88 | 0.96 | 0.89 | 1.01 |
| CollabLLM | 0.95 | 0.94 | 1.00 | 0.99 |
The main safety results are shown in Table 8, which presents the average harm scores across the four categories. Although all queries were adversarial and received high harm scores (typically between 4 and 7 on a 0–7 scale), both the Llama-3.1-8B baseline and CollabLLM produced responses that were, on average, very safe. Most scores are in the 0–1 range, which corresponds to “very low” harm. CollabLLM shows slightly lower harm in the Sexual and Hate categories and slightly higher harm in the other two. In terms of defect rate, CollabLLM produced only one response deemed unsafe by the SDK (out of 80 queries × 4 categories = 320 evaluations), resulting in a pass rate of 99.7%. Coincidentally, this is the same pass rate as Llama-3.1-8B, which also had one failed evaluation.
Overall, these results are encouraging. They suggest that CollabLLM’s training did not degrade the safety capabilities of the original LLM, even though no additional safety alignment was performed during CollabLLM’s training.
Appendix D Prompts
D.1 User Simulator
D.2 Prompt for Proactive Base
D.3 System Prompt
D.4 Interactivity Metric by LLM Judge
For the prompt template below, the ITR results reported in Table 1 use weights , , and , with the final score rescaled as , as all methods achieve an average ITR score above 2.5. Please use the same configuration to reproduce the results shown in Table 1. Note that the absolute values of , , and do not affect the overall conclusions. In our most recent codebase, we adopt , , and to eliminate the need for rescaling.
D.5 Helpfulness Reward by LLM Judge
Appendix E Question Template and Example on Abg-CoQA
@@
Appendix F User Study
F.1 User Study Platform
@@
F.2 Analysis: Divergence Between Simulated and Real Users
While user simulators were employed exclusively during training due to the large-scale conversation demands of our Multiturn-aware Reward computation, we provide a comparative analysis to study the divergence between user simulators and real users. We summarize key differences and similarities in communication patterns between real and simulated users below:
| Differences | Similarities |
| 1) Real users tend to use shorter, fragmented sentences with grammatical errors; simulators produce more complete and polished responses. | 1) Both exhibit iterative content development—progressively revealing information rather than specifying everything upfront. |
| 2) Real users often shift direction mid-conversation and introduce specific personal details (e.g., “eight dogs”); simulated users remain more predictable and generic. | 2) Both emphasize accessibility—frequently requesting simplifications, examples, and actionable guidance. |
| 3) Real users express emotion more bluntly (e.g., “that’s awful”) and use informal language, abbreviations, or incomplete thoughts; simulators respond in a more neutral and formal tone. | 3) Both articulate preferences about content structure or style, and provide feedback when expectations are met or unmet. |
Although our models were trained using simulated users, the user study demonstrates that they generalize effectively to real users. This supports the feasibility of simulator-based training for scalable optimization, while also revealing opportunities to enhance the realism and diversity of user simulators.